# On the stability, complexity, and distribution of similarity classes of the longest edge bisection process for triangles

### v1

@@ Abstract.
The Longest Edge Bisection (LEB) of a triangle is performed by joining the midpoint of its longest edge to the opposite vertex. Applying this procedure iteratively produces an infinite family of triangles. Surprisingly, a classical result of Adler (1983) shows that for any initial triangle, this infinite family falls into finitely many similarity classes.
While the set of classes is finite, we show that a far smaller, stable subset of “fat” triangles, called terminal quadruples, effectively dominates the final mesh structure. We prove the following asymptotic area distribution result: for every initial triangle, the portion of area occupied by terminal quadruples tends to one, with the convergence occurring at an exponential rate. In fact, we provide the precise distribution of triangles in every step. We introduce the bisection graph and use spectral methods to establish this result.
@@ 1.1. Background
The Longest Edge Bisection (LEB) process is appealing because it is easy to implement and generates simple triangulations through a straightforward edge-splitting procedure. The field was formally initiated by the seminal 1975 paper by Rosenberg and Stenger [RS75]. They defined the iterative process, where in each step the median to the longest edge of the triangle is added, thereby creating two “daughter” triangles.
The most critical result from Rosenberg and Stenger’s work is the proof of non-degeneracy, which guarantees the geometric quality of the triangles generated throughout the iterative process. They established that for any initial triangle $\Delta_{0}$ with smallest angle of $\alpha_{0}>0$, any angle $\theta$ in any triangle generated by the infinite $\mathrm{LEB}$ process is bounded from below by $\frac{\alpha_{0}}{2}$.
Beyond preventing angle degradation, a useful triangulation process must also ensure that the mesh resolution systematically increases. Explicitly, the maximum diameter over all simplices in the triangulation converges to zero. This property, known as convergence of the mesh diameter, was first formally proven for $\mathrm{LEB}$ in arbitrary dimensions in 1978 by Kearfott [KEA78]. This result was later sharpened by Stynes [STY79, STY80] and by Adler [ADL83], who provided improved bounds on the rate of convergence of the diameter for various classes of triangles.
A landmark result was provided by Adler in 1983, who proved the Finite Similarity Classes Theorem [ADL83]. This theorem states that the infinite family of triangles generated by the iterative bisection of a single initial triangle falls into a finite number of similarity classes. This surprising, much stronger, and more profound result, has deep implications for the stability of the method.
Adler’s theorem provides a crucial qualitative guarantee, but it opens the door to a deeper, more quantitative analysis of the $\mathrm{LEB}$ process as a dynamical system. While we know that the variety of shapes is finite, this does not describe the long-term behavior or the ultimate distribution of triangles. Several fundamental questions naturally arise: Does the process converge to a stable, periodic set of shapes? If so, what are the geometric properties of these terminal sets? At what rate does the total area of the mesh become dominated by these stable shapes? Furthermore, how does the number of distinct terminal sets depend on the geometry of the initial triangle? A comprehensive understanding of the stability and complexity of the $\mathrm{LEB}$ process requires answering these questions. A precise statement of these questions is given in the following §1.2, and answering them is the main contribution of this paper. Partial results in these directions appeared in [GGR07].
@@ 1.2. Preliminaries and main results
The 2014 paper [PP14] by Perdomo and Plaza provided a new proof of Adler’s theorem, employing the machinery of hyperbolic dynamics in a moduli space of triangle shapes, modeled in the hyperbolic upper half-plane $\mathbb{H}$, to analyze the bisection process as a dynamical system. Given a triangle $\Delta_{0}$, by
@@
Relying on this identification, we refer to a point $z\in D$ as a triangle. For such $z\in D$, the bisection by the median to the longest edge defines two function $L,R:D\to D$, where $L(z)$ is the left triangle and $R(z)$ is the right triangle. As explained in [PP14], the region $D$ is subdivided into $6$ subregions by the geodesics $\mathrm{Re}(z)=\frac{1}{4}$, $\left|{z}\right|=\frac{1}{2}$ and $\left|{z-\frac{1}{2}}\right|=\frac{1}{2}$ (see Figure 1), and the functions $L$ and $R$ are defined in a piecewise manner, by a fixed Möbius or anti-Möbius transformation on each one of these regions. The formulas and the geometric interpretation of these transformations are presented in Table 1 (compare [PP14, Figure 4]). An illustrative application presenting the geometry of the various transformations is available at https://danielkalmanovich.wixsite.com/website/leb. The reader is referred to [AHL78, §3] for more details on the action by Möbius transformations on the upper half-plane.
@@
As we will show, an exceptional property of the dynamics of the $\mathrm{LEB}$ process is that, regardless of the initial triangle, the process eventually produces predominantly triangles from a certain subregion of $\text{I}\cup\text{IV}$ in $D$, with notable geometric properties. Furthermore, the orbit of these triangles is a periodic set with at most $4$ elements. Generically [^1] See §2.1., these sets are of size exactly $4$, thus we call them terminal quadruples. This analysis immediately raises several questions that motivate this paper.
@@
To answer these questions, we introduce the bisection graph (see §3) of a triangle, and use spectral methods to analyze it. Furthermore, we adopt and refine the methods used in [PP14] to study the geometric properties of the orbit of $z$ under iterative $\mathrm{LEB}$, denoted by $\mathrm{LEB}(z)$. We also denote by $q(z)$ the number of terminal quadruples that arise from $z$, and by $l(z)=\left|{\mathrm{LEB}(z)}\right|$. Our main results are the following.
@@
In view of the theorem above, the area of any initial triangle $z$ gets filled with triangles that belong to terminal quadruples, with an exponential rate as the $\mathrm{LEB}$ process progresses. Thus, the number of triangle shapes that one sees is essentially $4q(z)$ and not $l(z)$. For instance, taking an acute isosceles triangle with apex angle $\alpha$ we have $l(z)=\Theta\left(\log\left(\frac{1}{\alpha}\right)\right)$ [GGR07, Lemma 2], while according to Theorem 2 below we have $q(z)=1$. In general, the following theorems shed light on the quantity $q(z)$, for $z$ in the different regions.

### v2

@@ Abstract.
The Longest Edge Bisection of a triangle is performed by joining the midpoint of its longest edge to the opposite vertex. Applying this procedure iteratively produces an infinite family of triangles. Surprisingly, a classical result of Stynes (1980) shows that for any initial triangle, this infinite family falls into finitely many similarity classes.
While the set of classes is finite, it turns out that a far smaller, stable subset of “fat” triangles effectively dominates the final mesh structure. We prove the following asymptotic area distribution result: for every initial triangle, the portion of area occupied by terminal quadruples tends to one, with the convergence occurring at an exponential rate. In fact, we provide the precise distribution of triangles in every step. We introduce the bisection graph and use spectral methods to establish this result.
@@ 1.1. Background
The Longest Edge Bisection (LEB) process is a straightforward iterative procedure formally initiated by Rosenberg and Stenger [RS75]. In each step, the median to the longest edge is added, splitting the triangle into two daughter triangles. Rosenberg and Stenger established the critical property of non-degeneracy, which guarantees geometric quality. They proved that for an initial triangle $\Delta_{0}$ with smallest angle $\alpha_{0}$, every angle generated by the infinite process is bounded from below by $\frac{\alpha_{0}}{2}$. Additionally, a useful triangulation process requires the convergence of the mesh diameter to zero. Kearfott [KEA78] first proved this for LEB in arbitrary dimensions; later, Stynes [STY79, STY80] and Adler [ADL83] provided sharpened bounds on the convergence rates.
A landmark result, known as the Finite Similarity Classes (FSC) Theorem, was established by Stynes [STY80, Corollary 1]. This theorem asserts that the infinite family of triangles generated by iteratively bisecting a single initial triangle falls into a finite number of similarity classes. This result was later reproved by Adler [ADL83] using an elegant inductive argument, and subsequently by Perdomo and Plaza [PP14], who recast the analysis within the framework of hyperbolic geometry, a setting that proved to be particularly natural.
In their paper, Rosenberg and Stenger identified a specific region where triangles enter a periodic orbit of length $4$ under the $\mathrm{LEB}$ process [RS75, Lemma 4]; we refer to these orbits as terminal sets. Building on this work and the FSC Theorem, Stynes [STY80, Corollary 2] demonstrated that the proportion of the area occupied by triangles in these terminal sets converges to $1$.
These results pave the way for a deeper, quantitative analysis of the $\mathrm{LEB}$ process as a dynamical system. While it is known that the variety of shapes is finite and that terminal sets asymptotically exhaust the area of the initial triangle, this does not describe the rate of convergence or the precise distribution of triangle shapes. Several fundamental questions naturally arise:
At what rate does the total area of the mesh become dominated by these stable shapes? What are the specific geometric properties of these terminal sets? Furthermore, how does the number of distinct terminal sets depend on the geometry of the initial triangle? Answering these questions is essential for a comprehensive understanding of the stability and complexity of the $\mathrm{LEB}$ process. We provide a precise formulation of these problems in §1.2 and address them as the main contribution of this paper. Partial results in these directions were previously noted in [GGR07].
@@ 1.2. Preliminaries and main results
The 2014 paper [PP14] by Perdomo and Plaza provided a new proof of the FSC theorem, employing the machinery of hyperbolic dynamics in a moduli space of triangle shapes, modeled in the hyperbolic upper half-plane $\mathbb{H}$, to analyze the bisection process as a dynamical system. Given a triangle $\Delta_{0}$, by
@@
Relying on this identification, we refer to a point $z\in D$ as a triangle. For such $z\in D$, the bisection by the median to the longest edge defines two function $L,R:D\to D$, where $L(z)$ is the left triangle and $R(z)$ is the right triangle. As explained in [PP14], the region $D$ is subdivided into $6$ subregions by the geodesics $\mathrm{Re}(z)=\frac{1}{4}$, $\left|{z}\right|=\frac{1}{2}$ and $\left|{z-\frac{1}{2}}\right|=\frac{1}{2}$ (see Figure 1), and the functions $L$ and $R$ are defined in a piecewise manner, by a Möbius or anti-Möbius transformation on each one of these regions. The formulas and the geometric interpretation of these transformations are presented in Table 1. An illustrative application presenting the geometry of the various transformations is available at https://danielkalmanovich.wixsite.com/website/leb. The reader is referred to [AHL78, §3] for more details on the action by Möbius transformations on the upper half-plane.
@@
As we will show, an exceptional property of the dynamics of the $\mathrm{LEB}$ process is that, regardless of the initial triangle, the process eventually produces predominantly triangles from a certain subregion of $\text{I}\cup\text{IV}$, with notable geometric properties. Furthermore, the orbit of such a triangle is periodic of size $4$, thus we call it a terminal quadruple. These facts immediately raise several questions that motivate this paper.
@@
To answer these questions, we introduce the bisection graph (see §3) of a triangle, and use spectral methods to analyze it. Furthermore, we adopt and refine the methods used in [PP14] to study the geometric properties of the orbit of $z$ under iterative $\mathrm{LEB}$. We denote by $\mathrm{LEB}(z)$ the orbit of $z$, by $q(z)$ the number of terminal quadruples that arise from $z$, and by $l(z)=\left|{\mathrm{LEB}(z)}\right|$. Our main results are the following.
@@
This theorem immediately implies Stynes’ result, stating that the area of any initial triangle $z$ gets filled with triangles that belong to terminal quadruples (see [STY80, Corollary 2]). Furthermore, Theorem 1 exhibits the precise distribution vector at every step and the rate of convergence.
It follows that the number of triangle shapes that one sees is essentially $4q(z)$ and not $l(z)$. For instance, taking an acute isosceles triangle with apex angle $\alpha$ we have $l(z)=\Theta\left(\log\left(\frac{1}{\alpha}\right)\right)$ [GGR07, Lemma 2], while according to Theorem 2 below we have $q(z)=1$. In general, the following theorems shed light on the quantity $q(z)$, for $z$ in the different regions.